Given position x = 2t + 5t2 (where x is in meters and t is in seconds):

A. Calculate the average velocity over the time interval t = 1 s to t = 4 s. Units: m.s-1

B. What is the instantaneous velocity at t = 4 s? Units: m.s-1

C. What is the acceleration of the object? Units: m.s-2

D. In which direction is the object accelerating?
Answer #1

x= 2t + 5t² V= dx - 2 + 10t at A V(1) = 2+ 10 = 12 m/s V (4) = 2+40 = 42 m/s 42 +12 Avaroge velocity + VC4) 2 27 m/s B) a du